OpenStudy (anonymous):

solve (x-h)^2+B^2=r^2......solve for B

OpenStudy (anonymous):

\[(x-h)^2+B^2=r^2\]

jimthompson5910 (jim_thompson5910):

start by subtracting \(\Large (x-h)^2\) from both sides

jimthompson5910 (jim_thompson5910):

after you isolate \(\Large B^2\) you can take the square root of both sides to isolate B itself.

OpenStudy (anonymous):

so i got \[B=\sqrt{r^2-x-h^2}\]

jimthompson5910 (jim_thompson5910):

it should be \[\Large B = \sqrt{r^2 - (x-h)^2}\] assuming B is nonnegative

OpenStudy (anonymous):

oh ok. can you help me out by testing on some inverse trig functions?

jimthompson5910 (jim_thompson5910):

post your question

OpenStudy (anonymous):

like give me questons similar to arcsin(tang1/2).

jimthompson5910 (jim_thompson5910):

ok how about `cos(arcsin(1/2))`

OpenStudy (anonymous):

ok. umm.. sqrt3/2?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

give me a few more. particularly where the range changes cause those confuse me.

jimthompson5910 (jim_thompson5910):

arcsin(cos(5pi/4))

OpenStudy (anonymous):

hmmm i have no clue.

jimthompson5910 (jim_thompson5910):

find the value of cos(5pi/4) first

OpenStudy (anonymous):

thats -sqrt2/2

jimthompson5910 (jim_thompson5910):

now find out what value of theta makes sin(theta) = -sqrt(2)/2 true keep in mind that the range of arcsine is -pi/2 <= y <= pi/2

OpenStudy (anonymous):

so would it be pi/4?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

-pi/4??

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

2 more please.

OpenStudy (anonymous):



OpenStudy (anonymous):

im trying to figure out sec(arcsin(x-1))

jimthompson5910 (jim_thompson5910):

what is b equal to in terms of x

OpenStudy (anonymous):

im writing an algebraic expression

OpenStudy (anonymous):

im solving for b.

OpenStudy (anonymous):

so (x-1)^2+B2=1....subtract (x-1)^2 on both sides.....then im left with B^2=1-(x-1)^2..... i square both sides..... now i have B= sqrt1-(x-1)^2

jimthompson5910 (jim_thompson5910):

yes \[\Large b = \sqrt{1-(x-1)^2}\]

OpenStudy (anonymous):

and now im suppose to relate that as a algebraic expression. so back to sec(arcsin(x-1))

OpenStudy (anonymous):

since im working inside out, sin is opp/hyp. so x-1/sqrt1-(x-1). but i dont know the secant.

OpenStudy (anonymous):

would it be 1/sqrt1-(x-1)^2???

jimthompson5910 (jim_thompson5910):

if you meant to say \[\Large \frac{1}{\sqrt{1-(x-1)^2}}\] then you are correct

OpenStudy (anonymous):

but my teacher does not like square roots as the denominator.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

then multiply top and bottom by \(\Large \sqrt{1-(x-1)^2}\) to rationalize the denominator

jimthompson5910 (jim_thompson5910):

doing that means \[\Large \frac{1}{\sqrt{1-(x-1)^2}}=\frac{\sqrt{1-(x-1)^2}}{1-(x-1)^2}\]

OpenStudy (anonymous):

All right. Thank you so much. I appreciate you taking the time to help me out!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Can you tell me what x+2)^2+B^2=1 is?

OpenStudy (anonymous):

i got sqrt1-(x-2)^2

jimthompson5910 (jim_thompson5910):

solving \[\Large (x+2)^2+B^2=1\] for B gives \[\Large B = \sqrt{1-(x+2)^2}\]

OpenStudy (anonymous):

all right thank you again!